OpenStudy (anonymous):

write an equation in the form y= mx+b for the line that has a slope of -5 and contains the point (7,0)

pooja195 (pooja195):

\[\huge~y-y_1=m(x-x_1)\] m=slope (x1,y1)

OpenStudy (rhystic):

So you have m in the equation which is the slope. thus you have y = -5x+ b now you need to solve for b. to do this take the y coordinate (0) and plug in to the equation and plug the x coordinate into the equation yielding -> 0 = -5(7) + b <- now solve for b. then you now can put it back into the form of y = mx + b

OpenStudy (anonymous):

so i multiply -5 from 7 will equal -35

OpenStudy (rhystic):

right. now solve for b .

OpenStudy (anonymous):

how do i solve for b

OpenStudy (rhystic):

so you found -35 right which is m*x in the equation y = mx + b you already have 0 for y meaning 0 = -35 + b you now need to isolate b which means b = what?

OpenStudy (anonymous):

b=-35

OpenStudy (rhystic):

not quite.

OpenStudy (rhystic):

 does that make sense?